OpenStudy (anonymous):

i need help with 1+1

OpenStudy (btaylor):

1 + 1 = 10 (if we are in binary)

OpenStudy (anonymous):

lie

OpenStudy (anonymous):

1+1=0 if we were in I.T

OpenStudy (anonymous):

1+1=2 [if we wher in america]

OpenStudy (anonymous):

ask u'r grade 1 teacher..i know he can solve that..

OpenStudy (anonymous):

can you count ?

OpenStudy (anonymous):

1+1 = 2

OpenStudy (anonymous):

not cool im out of school i graduated highschool

OpenStudy (anonymous):

if you can count, count the number of 1's in that expression

OpenStudy (anonymous):

that is your answer

OpenStudy (anonymous):

123456789

OpenStudy (anonymous):

i can count

OpenStudy (anonymous):

so how many 1's are in the expression?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

troll lolololololol

OpenStudy (anonymous):

you are right, good job :) ok problem solved